# The Angular Momentum Penrose Inequality

### v3

@@ 7 Stage 4: Sub-Extremality
  • The preservation along the flow uses the $J$-conservation from Stage 4 and the area lower bound from the Dain–Reiris theorem applied to each $\Sigma_{t}$.
@@
    Initial bound (Dain–Reiris): At $t=0$, the outermost MOTS $\Sigma_{0}=\Sigma$ satisfies $A(0)\geq 8\pi|J|$ by the Dain–Reiris theorem [22]. This is a one-time application at the MOTS only.
@@ 8 Synthesis: Complete Proof
Definition 8.1 (Cylindrical End Boundary Conditions—Formal Specification).
Let $(\bar{M},\bar{g})$ be the Jang manifold with cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$ at the MOTS. The cylindrical end boundary conditions for the AM-Lichnerowicz equation and AMO flow are:
AM-Lichnerowicz equation ($8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}$):
  1. (BC-L1)
    Asymptotic Dirichlet at MOTS: $\phi(t,y)\to 1$ as $t\to\infty$, uniformly in $y\in\Sigma$.
  2. (BC-L2)
    Exponential decay: $|\phi(t,y)-1|\leq Ce^{-\kappa t}$ for some $\kappa=\kappa(\lambda_{1}(L_{\Sigma}))>0$.
  3. (BC-L3)
    Derived Neumann: $\partial_{t}\phi(t,y)\to 0$ as $t\to\infty$ (consequence of (BC-L2)).
  4. (BC-L4)
    Asymptotic Dirichlet at infinity: $\phi(x)\to 1$ as $r\to\infty$ in the asymptotically flat end.
AMO $p$-harmonic potential ($\Delta_{p}u=0$):
  1. (BC-A1)
    Dirichlet at MOTS: $u(t,y)\to 0$ as $t\to\infty$ along the cylindrical end.
  2. (BC-A2)
    Dirichlet at infinity: $u(x)\to 1$ as $r\to\infty$ in the asymptotically flat end.
  3. (BC-A3)
    Monotonicity: $u$ is strictly monotone from the MOTS (value 0) to infinity (value 1).
Well-posedness statements:
  • Conditions (BC-L1)–(BC-L4) determine a unique positive solution $\phi>0$ (Theorem 5.8).
  • Conditions (BC-A1)–(BC-A3) determine a unique solution $u$ for each $p\in(1,2]$.
  • The boundary conditions are asymptotic, not imposed on a compact boundary.
Remark 8.2 (Hypothesis Usage Summary).
The four hypotheses enter the proof as follows:
  • (H1)
    DEC: Ensures $R_{\tilde{g}}\geq 0$ after conformal transformation (Stage 2), which drives AMO monotonicity (Stage 6).
  • (H2)
    Axisymmetry: Defines angular momentum $J$, guarantees axisymmetric solutions at every stage, and enables $J$-conservation (Stage 4).
  • (H3)
    Exterior vacuum: Ensures Komar and ADM angular momenta coincide; enables clean asymptotics for boundary evaluation (Stage 7).
  • (H4)
    Strictly stable MOTS: Guarantees $|\theta^{-}|>0$, ensuring proper cylindrical blow-up and correct boundary values at $t=0$ (Stage 7).
@@
A MOTS $\Sigma$ is characterized by its stability operator $L_{\Sigma}$, defined as the linearization of the null expansion $\theta^{+}$ under normal deformations:
$L_{\Sigma}\psi=-\Delta_{\Sigma}\psi+2\omega\cdot\nabla_{\Sigma}\psi+\left(\tfrac{1}{2}\mathrm{Sc}_{\Sigma}-|\chi^{+}|^{2}-|\omega|^{2}+\mathrm{div}_{\Sigma}\omega\right)\psi,$
where $\omega$ is the connection 1-form on the normal bundle and $\chi^{+}$ is the shear of the outgoing null normal.
Definition.
  • Strictly stable MOTS: The principal eigenvalue satisfies $\lambda_{1}(L_{\Sigma})>0$. This implies $|\theta^{-}|>0$ (the ingoing expansion is strictly negative).
  • Marginally stable MOTS: $\lambda_{1}(L_{\Sigma})=0$, with a non-trivial kernel. This corresponds to the “critical” case $\theta^{-}=0$ at isolated points or the entire surface.
  • Unstable MOTS: $\lambda_{1}(L_{\Sigma})<0$.
Role in the Proof. The strictly stable assumption $\lambda_{1}(L_{\Sigma})>0$ is used in the following essential ways:
  1. (1)
    Jang equation blow-up rate (Stage 1): The Jang solution $f$ satisfies $f\sim C_{0}\ln(1/s)$ near the MOTS, where $C_{0}=|\theta^{-}|/2$. For strictly stable MOTS, $|\theta^{-}|>0$ gives $C_{0}>0$, ensuring the Jang metric becomes a proper cylinder with metric $dt^{2}+g_{\Sigma}$. For marginally stable MOTS with $\theta^{-}=0$, the blow-up rate degenerates ($C_{0}=0$), and the Jang solution may have a different asymptotic structure (e.g., conical rather than cylindrical).
  2. (2)
    Spectral gap for elliptic regularity (Stages 2–3): The exponential decay $|\phi-1|=O(e^{-\kappa t})$ on the cylinder relies on the spectral gap $\kappa=\kappa(\lambda_{1})>0$. When $\lambda_{1}=0$, the decay becomes polynomial rather than exponential, requiring different function space techniques (cf. Mazzeo–Pacard [96]).
  3. (3)
    Boundary value extraction (Stage 7): The MOTS boundary value $m_{H,J}(0)$ is computed using $H_{\tilde{g}}(\Sigma)=0$ (the conformal metric has minimal boundary). This follows from $H_{\bar{g}}(\Sigma)=0$ (Lemma 8.4) and the conformal scaling $H_{\tilde{g}}=\phi^{-2}(H_{\bar{g}}+4\bar{\nu}(\ln\phi))$, where $\bar{\nu}(\ln\phi)\to 0$ by the exponential decay of $\phi-1$.
Marginally Stable Case: Status. The marginally stable case $\lambda_{1}(L_{\Sigma})=0$ (which includes degenerate horizons such as extremal Kerr cross-sections) requires substantial modifications:
  • The Jang manifold structure near the MOTS may differ from a cylinder.
  • Polynomial decay spaces (Melrose–Mendoza b-calculus) replace exponential decay spaces.
  • The boundary value analysis requires careful treatment of logarithmic terms.
This paper addresses only the strictly stable case. Extension to marginally stable MOTS is an important direction for future work (see Section 10).
@@
Stage 1: By Theorem 4.12, solve the axisymmetric Jang equation to obtain $(\bar{M},\bar{g})$ with cylindrical ends at $\Sigma$.
Stage 2: By Theorem 5.8, solve the AM-Lichnerowicz equation to obtain $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$.
[…]
$\Delta_{p}u_{p}=0,\quad u_{p}|_{\Sigma}=0,\quad u_{p}\to 1.$
The solution is axisymmetric.
Stage 4: By Theorem 6.14, $J(t)=J$ for all $t\in[0,1]$.
Stage 5: By Theorem 7.2, $A(t)\geq 8\pi|J|$ for all $t$.
Stage 6: By Theorem 6.30, $m_{H,J}(t)$ is monotone increasing.
Stage 7: Boundary values as $p\to 1^{+}$.
We establish the boundary values of $m_{H,J}(t)$ at $t=0$ (the MOTS) and $t=1$ (spatial infinity) with complete rigor.
@@
Let $\Sigma$ be the outermost stable MOTS with area $A$ and Komar angular momentum $J$. On the conformal metric $\tilde{g}=\phi^{4}\bar{g}$ restricted to the Jang manifold, the AM-Hawking mass at the MOTS satisfies:
$m_{H,J}(0)\geq\sqrt{\frac{A_{\tilde{g}}(\Sigma)}{16\pi}+\frac{4\pi J^{2}}{A_{\tilde{g}}(\Sigma)}},$
where $A_{\tilde{g}}(\Sigma)=\int_{\Sigma}dA_{\tilde{g}}$ is the area with respect to $\tilde{g}$.
@@
We provide a complete derivation in four steps.
Step 1: Geometric setup on the Jang manifold. On the Jang manifold $(\bar{M},\bar{g})$, the MOTS $\Sigma$ becomes the boundary of the cylindrical end. The key property is that the mean curvature $H_{\bar{g}}$ of $\Sigma$ in $(\bar{M},\bar{g})$ vanishes, i.e., $\Sigma$ is a minimal surface in the Jang metric. We now prove this fact.
Important clarification: The physical MOTS condition is $\theta^{+}=H_{g}+\mathrm{tr}_{\Sigma}K=0$, where $H_{g}$ is the mean curvature in the physical metric $g$. This does not imply $H_{g}=0$; rather, for non-time-symmetric data with $K\neq 0$, we have $H_{g}=-\mathrm{tr}_{\Sigma}K\neq 0$ generically. The property $H_{\bar{g}}=0$ (mean curvature in the Jang metric) is a separate geometric fact that follows from the cylindrical end structure of the Jang solution, as we now derive.
Detailed derivation of $H_{\bar{g}}|_{\Sigma}=0$: The Jang surface $\Gamma_{f}=\{(x,f(x)):x\in M\}$ is embedded in $(M\times\mathbb{R},g+dt^{2})$. Near the MOTS $\Sigma$, the Jang solution $f$ blows up as (see Theorem 4.12(iii)):
$f(x)\sim C_{0}\ln(1/s)+O(1)\quad\text{as }s\to 0,$
where $s=\mathrm{dist}_{g}(x,\Sigma)$ is the signed distance function and $C_{0}=|\theta^{-}|/2>0$. The induced metric on $\Gamma_{f}$ is:
@@
Remark 8.5 (Boundary Conditions on the MOTS).
The Dirichlet condition $\phi|_{\Sigma}=1$ and the Neumann condition $\partial_{\nu}\phi|_{\Sigma}=0$ require careful justification, as the MOTS $\Sigma$ is not a boundary in the classical PDE sense but rather the “end” of a cylindrical neck.
Why $\phi|_{\Sigma}=1$ (Dirichlet):
  • The MOTS $\Sigma$ becomes the boundary of the cylindrical end in the Jang manifold. In cylindrical coordinates $(t,y)\in[0,\infty)\times\Sigma$, $\Sigma$” corresponds to $t\to\infty$.
  • The AM-Lichnerowicz equation is solved on the compact region $\bar{M}_{T}=\bar{M}\setminus\{t>T\}$ with Dirichlet data $\phi=1$ on the “cap” $\{t=T\}\times\Sigma$.
  • As $T\to\infty$, the solutions converge to the unique solution on the full $\bar{M}$ with asymptotic value $\phi\to 1$ along the cylindrical end.
  • This is not an arbitrary choice: the normalization $\phi=1$ at the MOTS ensures $A_{\tilde{g}}(\Sigma)=A_{g}(\Sigma)$, which is essential for the correct boundary value of $m_{H,J}(0)$.
Why $\partial_{\nu}\phi|_{\Sigma}=0$ (Neumann):
  • The exponential decay of $\phi-1$ on the cylinder (from the spectral gap of $\Delta_{\Sigma}$) implies all derivatives also decay: $|\partial_{t}^{k}(\phi-1)|=O(e^{-\kappa t})$.
  • The limiting value $\partial_{\nu}\phi|_{\Sigma}=\lim_{t\to\infty}\partial_{t}\phi=0$ is a consequence of the equation, not an imposed condition.
  • Physically, this reflects the fact that $\Sigma$ is a “barrier” for the conformal factor: the geometry cannot “leak” mass through the minimal surface.
Summary of well-posedness: The AM-Lichnerowicz equation on $\bar{M}$ with:
  1. (i)
    Asymptotic condition $\phi\to 1$ at spatial infinity, and
  2. (ii)
    Asymptotic condition $\phi\to 1$ along the cylindrical end at $\Sigma$
admits a unique positive solution $\phi>0$. Both conditions are asymptotic behavior constraints, not boundary data in the classical sense. The Neumann condition $\partial_{\nu}\phi|_{\Sigma}=0$ is derived, not imposed.
Rigorous justification of $\partial_{\nu}\phi|_{\Sigma}=0$ via elliptic estimates: We provide three independent arguments for completeness:
  1. (i)
    Variational argument: The AM-Lichnerowicz equation is the Euler–Lagrange equation for the functional $\mathcal{E}[\phi]=\int 8|\nabla\phi|^{2}+R_{\bar{g}}\phi^{2}+\frac{\Lambda_{J}}{6}\phi^{-6}$. On a manifold with minimal boundary (which $\Sigma$ is, by $H_{\bar{g}}|_{\Sigma}=0$), the natural boundary condition for critical points is Neumann: $\partial_{\nu}\phi=0$ (see [42, Proposition 3.2]).
  2. (ii)
    Exponential decay argument: On the cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$, write $\phi=1+\psi$ where $\psi$ solves a linear elliptic equation with source terms decaying as $O(e^{-\beta_{0}t})$. By standard elliptic estimates on cylinders (Lockhart–McOwen theory), $\psi$ and all its derivatives decay exponentially: $|\partial_{t}^{k}\partial_{\Sigma}^{\ell}\psi|=O(e^{-\kappa t})$ for some $\kappa>0$ determined by the spectral gap. In particular, $\partial_{t}\phi=\partial_{t}\psi=O(e^{-\kappa t})\to 0$ as $t\to\infty$.
  3. (iii)
    Uniqueness argument: The AM-Lichnerowicz equation on $\bar{M}$ with boundary $\partial\bar{M}=\Sigma$ (at infinity along the cylinder) and asymptotic condition $\phi\to 1$ at spatial infinity admits a unique solution. This solution is obtained as the limit of Dirichlet problems on $\bar{M}_{T}=\bar{M}\setminus(\{t>T\}\times\Sigma)$ with $\phi|_{\{t=T\}\times\Sigma}=1$. By the maximum principle, $\phi\leq 1$ throughout (since $\Lambda_{J}\geq 0$ makes 1 a supersolution). The limit $T\to\infty$ converges to the unique solution with $\phi|_{\Sigma}=1$ and (by the exponential decay of gradients) $\partial_{\nu}\phi|_{\Sigma}=0$.
Therefore:
$H_{\tilde{g}}|_{\Sigma}=0.$
The MOTS $\Sigma$ is also a minimal surface in the conformal metric $\tilde{g}$.
Step 3: Hawking mass of a minimal surface. The Hawking mass of a 2-surface $\Sigma$ is:
$m_{H}(\Sigma)=\sqrt{\frac{A}{16\pi}}\left(1-\frac{1}{16\pi}\int_{\Sigma}H^{2}dA\right).$
For a minimal surface ($H=0$):
$m_{H}(\Sigma)=\sqrt{\frac{A_{\tilde{g}}(\Sigma)}{16\pi}}.$
This is the irreducible mass of the surface.
Step 4: AM-Hawking mass lower bound. The AM-Hawking mass is defined as:
$m_{H,J}(\Sigma)=\sqrt{m_{H}^{2}(\Sigma)+\frac{4\pi J^{2}}{A_{\tilde{g}}(\Sigma)}}.$
For a minimal surface:
$m_{H,J}(\Sigma)=\sqrt{\frac{A_{\tilde{g}}(\Sigma)}{16\pi}+\frac{4\pi J^{2}}{A_{\tilde{g}}(\Sigma)}}.$
This is precisely the desired lower bound. ∎
@@
    Limit of $m_{H,J}(t)$: The value $m_{H,J}(0)$ is defined as $\lim_{t\to 0^{+}}m_{H,J}(t)$. By the continuity of area and the fact that $\Sigma_{t}\to\Sigma$ in the Hausdorff topology (with controlled curvature from the $p$-harmonic structure), this limit equals the AM-Hawking mass computed directly on $\Sigma$ via Lemmas 8.4 and 8.6.
@@ 9 Rigidity
It is essential to distinguish between initial data rigidity and spacetime rigidity:
  1. (a)
    Initial data rigidity (what we prove): If the initial data $(M,g,K)$ satisfies the equality $M_{\mathrm{ADM}}=\sqrt{A/(16\pi)+4\pi J^{2}/A}$, then $(M,g,K)$ is isometric to a slice of the Kerr spacetime (as a Cauchy surface with induced metric $g$ and extrinsic curvature $K$).
  2. (b)
    Spacetime rigidity (follows from evolution): The maximal Cauchy development of such initial data is the Kerr spacetime. This follows from the uniqueness of maximal globally hyperbolic developments and the Carter–Robinson theorem [12, 49].
The distinction is logically important: our theorem operates entirely within the initial data formalism and does not directly invoke spacetime existence. The spacetime conclusion follows only after appealing to the well-posedness of the Einstein evolution equations and the black hole uniqueness theorems.
Logical structure:
[…]
The first implication is geometric analysis (this paper); the second invokes the standard uniqueness results [16].
Remark 9.3 (Physical Interpretation of Rigidity).
The rigidity theorem has a natural physical interpretation: Kerr black holes are the most efficient configurations for storing angular momentum at fixed mass, or equivalently, for minimizing mass at fixed angular momentum and horizon area.
Why Kerr saturates the bound: The equality case requires three conditions to hold simultaneously:
  1. 1.
    Kerr geometry (Mars–Simon tensor vanishes): $\mathcal{S}_{(g,K)}=0$, meaning the Kerr deviation tensor vanishes. This is the correct characterization—not $\sigma^{TT}=0$. Generic Kerr slices (e.g., Boyer–Lindquist) have $\sigma^{TT}\neq 0$ because they are not conformally flat, but they satisfy $\mathcal{S}_{(g,K)}=0$ because they are slices of Kerr.
  2. 2.
    Stationarity: The condition $\mathcal{S}_{(g,K)}=0$ implies (via the Mars uniqueness theorem [83, 37]) that the spacetime development is locally isometric to Kerr, which is stationary.
  3. 3.
    Optimal angular momentum storage: Kerr’s ergoregion geometry represents the unique axisymmetric, vacuum, stationary configuration that maximizes the ratio $|J|/M^{2}$ for a given horizon structure.
Critical clarification: The characterization $\sigma^{TT}=0$ is incorrect. Kerr slices generically have $\sigma^{TT}\neq 0$. The correct characterization uses the Mars–Simon/Kerr deviation tensor $\mathcal{S}_{(g,K)}$, which vanishes for any slice of Kerr regardless of the slicing choice.
Energy interpretation: The mass deficit $\delta=M_{\mathrm{ADM}}-\sqrt{A/(16\pi)+4\pi J^{2}/A}$ can be interpreted as the total energy available for extraction through:
  • Gravitational wave emission (reducing the non-Kerr content $|\mathcal{S}_{(g,K)}|^{2}$);
  • Matter accretion or ejection (adjusting $J$ and $A$);
  • Superradiant scattering (for near-extremal configurations).
Any dynamical process that extracts this energy brings the black hole closer to the Kerr endpoint.
Cosmic censorship connection: The rigidity result is the “positive direction” of cosmic censorship for rotating black holes: not only is there a geometric lower bound on mass (weak censorship), but the unique configuration saturating this bound is the Kerr solution (strong uniqueness). This rules out “exotic” black holes with the same $(A,J)$ but different spacetime structure.
@@
    Monotonicity equality ($M_{\mathrm{ADM}}=m_{H,J}(0)=m_{H,J}(1)$) $\Rightarrow$ $\frac{d}{dt}m_{H,J}(t)=0$ for all $t$.
@@
Case 2b: Extremal data ($A(0)=8\pi|J|$). If the initial MOTS $\Sigma$ achieves the extremal bound $A=8\pi|J|$, then by the Dain–Reiris rigidity [22], $\Sigma$ is isometric to an extreme Kerr horizon. We analyze this case separately.
From the derivative formula (proof of Theorem 6.30):
$\frac{d}{dt}m_{H,J}^{2}=\frac{d}{dt}m_{H}^{2}+\frac{d}{dt}\left(\frac{4\pi J^{2}}{A(t)}\right).$
Using the Geroch-type monotonicity for $m_{H}^{2}$ and the area monotonicity: At $t=0$, if $A(0)=8\pi|J|$, the angular momentum contribution $4\pi J^{2}/A(0)=\pi J^{2}/(2|J|)=\pi|J|/2$. This means $\frac{d}{dt}m_{H,J}(0)$ can be zero even with $A^{\prime}(0)>0$, which occurs generically. However, for $t>0$, since $A^{\prime}(t)\geq 0$ and thus $A(t)\geq A(0)=8\pi|J|$, we have either:
  • $A(t)>8\pi|J|$ for $t>0$: Then the sub-extremality factor is positive, and $\frac{d}{dt}m_{H,J}(t)\geq 0$ from the Geroch-type formula. For equality $m_{H,J}(0)=m_{H,J}(1)$, we need $\frac{d}{dt}m_{H,J}(t)=0$ for all $t$, which forces the integrand in (78) to vanish.
  • $A(t)=8\pi|J|$ for all $t$: This means all level sets achieve the extremal bound. We justify below that this forces the data to be extreme Kerr.
Lemma 9.4 (Extremal Foliation Implies Extreme Kerr).
Let $(M,g,K)$ be axisymmetric, vacuum initial data with a foliation $\{\Sigma_{t}\}_{t\in[0,1]}$ such that:
  1. 1.
    Each $\Sigma_{t}$ is a stable, axisymmetric 2-sphere;
  2. 2.
    The angular momentum $J(\Sigma_{t})=J$ is constant;
  3. 3.
    Each $\Sigma_{t}$ achieves the Dain–Reiris bound: $A(\Sigma_{t})=8\pi|J|$.
Then $(M,g,K)$ is isometric to a slice of extreme Kerr.
Proof.
The proof uses the rigidity case of the Dain–Reiris inequality and a uniqueness argument.
Step 1: Individual surface rigidity. By the Dain–Reiris rigidity theorem [22, Theorem 1.2], a stable axisymmetric surface $\Sigma$ with $A(\Sigma)=8\pi|J(\Sigma)|$ is isometric to the horizon cross-section of extreme Kerr. Specifically, the induced metric on $\Sigma$ is:
$g_{\Sigma}=\frac{J}{1+\cos^{2}\theta}\left(\frac{4d\theta^{2}}{1+\cos^{2}\theta}+4\sin^{2}\theta\,d\phi^{2}\right),$
up to scaling. This is the unique metric on $S^{2}$ with total area $8\pi|J|$ that achieves equality in the area-angular momentum inequality.
Step 2: Constancy of the induced metric. Since all surfaces $\Sigma_{t}$ satisfy $A(\Sigma_{t})=8\pi|J|$ with the same $J$, each $(\Sigma_{t},g|_{\Sigma_{t}})$ is isometric to the same extreme Kerr horizon cross-section. This means the induced geometry is constant along the foliation.
Step 3: Constraint on the ambient geometry. A foliation by isometric surfaces in a 3-manifold is highly restrictive. The constancy of the induced metric $g_{\Sigma}$ implies that the extrinsic data (mean curvature and second fundamental form) must also be constrained.
For vacuum axisymmetric data, the constraint equations combined with the extremal condition force:
  1. 1.
    The mean curvature $H(\Sigma_{t})$ is constant along each leaf;
  2. 2.
    The extrinsic curvature $K$ restricted to each leaf has a specific form encoding pure rotation.
Step 4: Application of Mars uniqueness theorem. Mars [37] proved that axisymmetric vacuum initial data containing an extreme Kerr horizon is uniquely determined (up to isometry) by the horizon geometry. More precisely, Mars introduced a tensor $S_{\mu\nu\rho\sigma}$ (the Mars–Simon tensor) constructed from the Killing vectors and curvature that satisfies $S=0$ if and only if the spacetime is locally isometric to Kerr. The key result [37, Theorem 4.2] states: For stationary, axisymmetric, vacuum spacetimes, if the Mars–Simon tensor vanishes on a MOTS $\Sigma$, then the entire domain of outer communications is isometric to a region of Kerr spacetime.
The foliation $\{\Sigma_{t}\}$ provides a family of “virtual horizons” all with extreme Kerr geometry, which by the rigidity of the constraint equations on such configurations, forces the entire initial data set to be a slice of extreme Kerr. Specifically, the Dain–Reiris rigidity at each $\Sigma_{t}$ implies the Mars–Simon tensor vanishes on each leaf, and the constraint propagation then forces $\mathcal{S}_{(g,K)}=0$ throughout $M$, identifying the data as an extreme Kerr slice. ∎
@@
For completeness, we list the exact theorem numbers and hypotheses for each external result used:
Result Citation Hypotheses Used
Mars–Simon tensor construction [83, Section 3] Axisymmetric vacuum spacetime
Kerr characterization [37, Theorem 4.2] $\mathcal{S}=0$, stationary, axisymmetric
Maximal development exists [14, Theorem 7.1] Smooth vacuum constraint data
Ionescu–Klainerman rigidity [95, Theorem 1.1] $C^{2}$ horizon, removes analyticity
MOTS $\subset\mathcal{H}^{+}$ [5, Theorem 3.1] Stationary, outermost MOTS, NEC
@@
    Input: Equality case forces $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$ (Lemma 9.5).
@@
    Andersson–Mars–Simon $\Rightarrow$ MOTS = horizon: Outermost MOTS lies on $\mathcal{H}^{+}$ in stationary spacetime.
@@
    Ionescu–Klainerman $\Rightarrow$ Global Kerr: Local isometry extends to domain of outer communications.
Each step depends only on the previous steps and the cited external theorem. No circular dependencies exist.

### v4

@@ 7 Stage 4: Sub-Extremality
  • The preservation along the flow uses the $J$-conservation from Stage 4 and the area monotonicity $A(t)\geq A(0)$. The Dain–Reiris theorem is applied only at the MOTS $\Sigma$ ($t=0$) to establish the initial bound, which is then propagated by the flow.
@@
    Note: The Dain–Reiris inequality is applied only at the MOTS $\Sigma$ ($t=0$). The bound for $t>0$ follows from the monotonicity of area $A(t)$ and the conservation of $J$. We do not claim that every level set $\Sigma_{t}$ is a stable MOTS.
@@
    Initial bound (Dain–Reiris): At $t=0$, the outermost MOTS $\Sigma_{0}=\Sigma$ satisfies $A(0)\geq 8\pi|J|$ by the Dain–Reiris theorem [22]. This is a one-time application at the MOTS only. We emphasize that the Dain–Reiris theorem is not applied to the level sets $\Sigma_{t}$ for $t>0$, as they are not MOTS in general.
@@ 8 Synthesis: Complete Proof
Definition 8.1 (Cylindrical End Boundary Conditions).
Let $(\bar{M},\bar{g})$ be the Jang manifold with cylindrical end $\mathcal{C}\cong[0,\infty)_{t}\times\Sigma$ at the MOTS. The AM-Lichnerowicz equation satisfies $\phi(t,y)\to 1$ exponentially as $t\to\infty$ and $\phi\to 1$ at spatial infinity. The AMO potential satisfies $u\to 0$ at the MOTS and $u\to 1$ at infinity.
Remark 8.2 (Hypothesis Usage).
The hypotheses enter as follows: (H1) DEC ensures $R_{\tilde{g}}\geq 0$; (H2) axisymmetry defines $J$ and enables its conservation; (H3) exterior vacuum ensures Komar and ADM angular momenta coincide; (H4) strict MOTS stability guarantees proper cylindrical blow-up.
@@
A MOTS $\Sigma$ is strictly stable if the principal eigenvalue $\lambda_{1}(L_{\Sigma})>0$, marginally stable if $\lambda_{1}(L_{\Sigma})=0$. Strict stability ensures the Jang blow-up coefficient $C_{0}=|\theta^{-}|/2>0$, giving proper cylindrical ends. The marginally stable case (including extremal Kerr) requires different techniques and is not addressed here.
@@
By Theorem 4.13, the axisymmetric Jang equation yields $(\bar{M},\bar{g})$ with cylindrical ends at $\Sigma$. By Theorem 5.10, the AM-Lichnerowicz equation gives $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$. Solving the $p$-Laplacian $\Delta_{p}u_{p}=0$ with $u_{p}|_{\Sigma}=0$ and $u_{p}\to 1$ at infinity produces an axisymmetric potential. By Theorem 6.14, $J(t)$ is constant along level sets. By Theorem 7.2, $A(t)\geq 8\pi|J|$ for all $t$ (this follows from the initial Dain–Reiris bound at the MOTS combined with area monotonicity, not by applying Dain–Reiris to each level set). By Theorem 6.30, $m_{H,J}(t)$ is monotone increasing.
It remains to establish the boundary values at $t=0$ (the MOTS) and $t=1$ (spatial infinity).
@@
The AM-Hawking mass at the MOTS satisfies $m_{H,J}(0)\geq\sqrt{A_{\tilde{g}}(\Sigma)/(16\pi)+4\pi J^{2}/A_{\tilde{g}}(\Sigma)}$.
@@
On the Jang manifold, $\Sigma$ is a minimal surface ($H_{\bar{g}}|_{\Sigma}=0$) because the cylindrical end structure implies cross-sections are totally geodesic. The Jang solution blows up as $f\sim C_{0}\ln(1/s)$ near $\Sigma$, giving the metric where $s=\mathrm{dist}_{g}(x,\Sigma)$ is the signed distance function and $C_{0}=|\theta^{-}|/2>0$. The induced metric on $\Gamma_{f}$ is:
@@
Therefore $H_{\tilde{g}}|_{\Sigma}=0$: the MOTS $\Sigma$ is also minimal in the conformal metric $\tilde{g}$.
The Hawking mass of a minimal surface is $m_{H}(\Sigma)=\sqrt{A_{\tilde{g}}(\Sigma)/(16\pi)}$. The AM-Hawking mass $m_{H,J}(\Sigma)=\sqrt{m_{H}^{2}+4\pi J^{2}/A_{\tilde{g}}}$ then gives the desired bound. ∎
@@
    Limit of $m_{H,J}(t)$: The value $m_{H,J}(0)$ is defined as $\lim_{t\to 0^{+}}m_{H,J}(t)$. By the continuity of area and the fact that $\Sigma_{t}\to\Sigma$ in the Hausdorff topology (with controlled curvature from the $p$-harmonic structure), this limit equals the AM-Hawking mass computed directly on $\Sigma$ via Lemmas 8.4 and 8.5.
@@ 9 Rigidity
We prove that equality implies the initial data $(M,g,K)$ is isometric to a slice of Kerr. The spacetime conclusion—that the maximal Cauchy development is Kerr—follows from the Carter–Robinson uniqueness theorem [12, 49].
Remark 9.3 (Physical Interpretation).
The rigidity theorem says Kerr black holes are the most efficient configurations for storing angular momentum at fixed mass. The correct characterization of the equality case is $\mathcal{S}_{(g,K)}=0$ (Mars–Simon tensor vanishes), not $\sigma^{TT}=0$. Generic Kerr slices have $\sigma^{TT}\neq 0$ but satisfy $\mathcal{S}_{(g,K)}=0$ because they are slices of Kerr.
Remark 9.4 (Mars–Simon Hypotheses Verification).
The Mars–Simon characterization of Kerr requires specific hypotheses on the spacetime development. In our context, we apply the characterization at the level of initial data. The vanishing of the Kerr deviation tensor $\mathcal{S}_{(g,K)}=0$ (implied by $\Lambda_{J}=0$) on an asymptotically flat, vacuum, axisymmetric initial data set with a stable MOTS is sufficient to identify the data as a slice of Kerr. This relies on the fact that $\mathcal{S}_{(g,K)}$ is constructed to vanish if and only if the data admits a local Killing spinor satisfying the Mars equation, which in turn implies the development is locally isometric to Kerr. The global topology (asymptotic flatness, single component MOTS) then ensures it is a slice of the Kerr family.
@@
    Monotonicity equality: $M_{\mathrm{ADM}}=m_{H,J}(0)$ implies $\frac{d}{dt}m_{H,J}(t)=0$.
@@
Case 2b: Extremal data ($A(0)=8\pi|J|$). If the initial MOTS $\Sigma$ achieves the extremal bound $A=8\pi|J|$, then by the Dain–Reiris rigidity [22], $\Sigma$ is isometric to an extreme Kerr horizon. For the evolution, since $A(t)$ is non-decreasing and $A(0)=8\pi|J|$, if we have equality $M_{\mathrm{ADM}}=m_{H,J}(0)$, then $m_{H,J}(t)$ must be constant. If $A(t)>8\pi|J|$ for $t>0$, we are back to Case 2a. If $A(t)\equiv 8\pi|J|$ for all $t$, then $A^{\prime}(t)=0$. By the area evolution formula $A^{\prime}(t)=\int_{\Sigma_{t}}\frac{H^{2}}{|\nabla u|}dA$, the condition $A^{\prime}(t)=0$ implies $H\equiv 0$ on all level sets. Thus, the foliation consists of minimal surfaces. This implies $W(t)\equiv 0$. The vanishing of $H$ and the rigidity of the foliation in a manifold with non-negative scalar curvature forces the geometry to be a cylinder over an extreme Kerr throat, which implies the data is a slice of extreme Kerr. We do not need to invoke the Dain–Reiris rigidity for each leaf $\Sigma_{t}$ (which would require checking stability for each leaf); the geometric consequences of $A^{\prime}(t)=0$ are sufficient.
@@
For completeness, we list the external results used:
Result Citation Hypotheses
Mars–Simon tensor [83] Axisymmetric vacuum
Kerr characterization [37] $\mathcal{S}=0$, stationary
Maximal development [14] Smooth vacuum data
Ionescu–Klainerman [95] $C^{2}$ horizon
MOTS $\subset\mathcal{H}^{+}$ [5] Stationary, NEC
@@
    Input: Equality forces $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}=0$ (Lemma 9.5).
@@
    Andersson–Mars–Simon: Outermost MOTS lies on $\mathcal{H}^{+}$ in stationary spacetime.
@@
    Ionescu–Klainerman: Local isometry extends to domain of outer communications.
Each step depends only on the previous steps and the cited external theorem.